OpenStudy (anonymous):

A countrys population in 1993 was 83 million. In 2001 it was 87 million. Estimate the population in 2004 using the exponential growth formula. Round your answer to the nearest million. P=Ae^kt

OpenStudy (anonymous):

find the rate k at 1993, t = 0 at 2001, t = 8 so in 2001, 87 million = (83 million) e^(8k) 87/83 = e^(8k) ln(87/83) = 8k \[\huge \frac{\ln(\frac{87}{83})}{8} = k\] find k then at 2004, t = 11

OpenStudy (anonymous):

I still don't understand

OpenStudy (anonymous):

sorry for the english P=Ae^(kt) P = population after t years A = population in 1993, 83 million k = rate of growth t = years after 1993 first we need to find k by substitute the data given for year 2001, shown in my first reply after you got k, just substitute t = 11 (year 2004) P = (83 million)e^(11k)